How much will be in someone's account at the end of 3 years if they deposit $700 and the simple interest is 5% per year?

Answer by stanbon(75887)

How much will be in someone's account at the end of 3 years if they deposit $700 and the simple interest is 5% per year?
---------------
A(t) = P(1 + rt)
-----
A(3) = 700(1 + 0.05*3) = 700*1.15 = 805
===========================================
Cheers,
Stan H.
==========